Evaluate the integral ∫x^3 sin⁡〖x〗dx


1
Expert's answer
2020-10-25T18:42:58-0400

"\int x^3sinx\, dx"



"Using\ the\ formula, \int Udv = UV -\int Vdu"


"\begin{aligned}

U = x^3 \qquad &; \qquad dv=sinx\,dx \\
\frac{du}{dx}=3x^2\qquad &; \qquad \int dv = \int sinx\, dx\\
du= 3x^2\,dx\qquad &; \qquad V = -cosx

\end{aligned}"


"\therefore \int x^3sinx\, dx = (x^3)(-cosx) - \int(-cosx)(3x^2\,dx)"


"\int x^3sinx\, dx = -x^3\,cosx - \int( -3x^2cosx\,dx)"


"\int x^3sinx\, dx = -x^3\,cosx - (-3)\int x^2cosx\,dx"


"\int x^3sinx\, dx = -x^3\,cosx +3 \int x^2cosx\,dx\ --(i)"




"\int x^2cosx\,dx ="


"\begin{aligned}

U = x^2\qquad &; \qquad dv = cosx\,dx\\
\frac{du}{dx}= 2x\qquad &; \qquad \int dv = \int cosx\,dx\\
du = 2x\,dx\qquad &; \qquad V = sinx

\end{aligned}"


"\therefore \int x^2cosx\, dx = (x^2)(sinx) - \int(sinx)(2x\,dx)"


"\int x^2cosx\, dx = x^2\,sinx - 2\int xsinx\,dx\ ---(ii)"




"\int xsinx\,dx ="


"\begin{aligned}

U = x\qquad &; \qquad dv = sinx\,dx\\
\frac{du}{dx}= 1\qquad &; \qquad \int dv = \int sinx\,dx\\
du = dx\qquad &; \qquad V = -cosx

\end{aligned}"


"\therefore \int xsinx\, dx = (x)(-cosx) - \int(-cosx)(dx)"


"\int xsinx\, dx = -xcosx + \int cosx\,dx"


"\int xsinx\, dx = -xcosx + sinx"


Substituting the value of "\int xsinx\,dx" into equation (ii),


"\int x^2cosx\, dx = x^2\,sinx - 2\int xsinx\,dx"


"\int x^2cosx\, dx = x^2\,sinx - 2(-xcosx+sinx)"


"\int x^2cosx\, dx = x^2\,sinx +2xcosx - 2sinx"



Substituting the value of "\int x^2cosx\,dx" into equation (i), we have


"\int x^3sinx\, dx = -x^3\,cosx +3 \int x^2cosx\,dx"


"\int x^3sinx\, dx = -x^3\,cosx +3 (x^2\,sinx +2xcosx - 2sinx)"


"\int x^3sinx\, dx = -x^3\,cosx +3x^2\,sinx +6xcosx - 6sinx + C"


"\therefore The\ integral\ of\ x^3sinx = -x^3\,cosx +3x^2\,sinx +6xcosx - 6sinx + C"